# ComSD: Balancing Behavioral Quality and Diversity in Unsupervised Skill Discovery

### v1

@@ 5 Conclusion
In this paper, we propose a novel unsupervised skill discovery method named ComSD. It improves both behavioral quality and diversity through contrastive learning from a multi-objective perspective. For numerical evaluation, we conduct comprehensive comparisons on different kinds of downstream adaptation tasks, considering behavioral diversity ignored by recent works. The superiority of ComSD over all other baselines is well verified by our employed evaluation. We hope (i) our ComSD can inspire more attention to intrinsic reward weighting algorithms, and (ii) our experimental settings can make a difference for unsupervised skill evaluation in future works.
@@ A.3 Skill Finetuning Evaluation
Appendix B Baseline Details
@@
(Eysenbach et al., 2018) is one of the most classical and original unsupervised skill discovery algorithms, trying to maximize the MI between skills and states. It employs the first MI decomposition, Eq. 1. It uses a discrete uniform prior distribution to guarantee the maximization of skill entropy $H(z)$. The negative state-conditioned entropy $-H(z|s)$ is estimated by a trainable discriminator ${\rm log}p(z|s)$ which computes the intrinsic reward. As a foundational work, it provides several reasonable evaluations of skill adaptation, of which skill finetuning and skill combination are employed in our experiments.
@@
(Lee et al., 2019) aims to learn a policy for which the state marginal distribution matches a given target state distribution. It optimizes the objective by reducing it to a two-player, zero-sum game between a state density model and a parametric policy. Like DIAYN, it is also based on the first decomposition (Eq. 1) of MI and employs discriminator training. The difference is that SMM explicitly maximizes the state entropy with intrinsic reward, which inspires lots of recent advanced works and our ComSD.
@@
(Liu & Abbeel, 2021a) first employs the second MI decomposition Eq. 2 for a better MI estimation. For state entropy estimation, it employs a popular particle-based entropy estimation proposed by APT (Liu & Abbeel, 2021b), which is proven effective and supports many advanced works (Laskin et al., 2022b; Yarats et al., 2021b) and our ComSD. For skill-conditioned entropy, it chooses the successor feature (Hansen et al., 2019), introducing it into the final intrinsic reward for an explicit maximization. The weight between different entropy estimations is fixed in APS. APS can’t guarantee the behavioral quality (exploration) well. Different from APS, our ComSD employs contrastive learning for better conditioned entropy estimation and designs a novel dynamic weighting algorithm (SMW) to overcome the exploration drop brought by explicit conditioned entropy maximization.
@@
(Laskin et al., 2022b) is a state-of-the-art robot behavior discovery method. It first introduces contrastive learning (Chen et al., 2020) into unsupervised skill discovery. It chooses the second MI decomposition, Eq. 2 with APT particle-based estimation for state entropy, like APS. The contrastive learning between state transitions and skill vectors is conducted for implicit skill-conditioned entropy maximization. The encoder learned by contrastive learning is further used for APT reward improvement. The behaviors produced by CIC are of high activity but not distinguishable. Different from CIC, we employ the contrastive results as diversity intrinsic rewards for explicit conditioned entropy maximization to improve behavioral diversity, with SMW to balance two entropy estimations for exploratory ability maintenance.
@@
(Yang et al., 2023) is another state-of-the-art method on URLB (Laskin et al., 2021) and 2D exploration (Campos et al., 2020). It tries to mitigate the exploitation problem in CIC by a novel MI objective, $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill. It provides theoretical proof to show that their novel MI objective serves as the upper bound of the previous MI objective. However, BeCL can’t generate enough dynamic robot behaviors, and their intrinsic reward computational consumption is also much larger than other approaches.
@@ Appendix D Additional Analysis
D.1 What Skills Do ComSD and Competitive Baselines Discover? & Why does ComSD exhibit better adaptation performance than others?
Figure 6: Visualization for representative behaviors discovered by CIC and APS. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4. Skill AKD ranges from 7 to 10 for CIC and 0 to 2 for APS, which means they can’t discover qualified behaviors at different activity levels.
Figure 7: Visualization for representative behaviors discovered by ComSD. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4. Skill AKD ranges from 0 to 10 for ComSD. The results demonstrate that our ComSD can produce a qualified skill set consisting of diverse behaviors at different activity levels, which recent advanced methods cannot.
We provide the skill visualization and corresponding skill AKD of the two most competitive baselines, APS (Liu & Abbeel, 2021a) and CIC (Laskin et al., 2022b), in Figure 6. The visualization and skill AKD of our ComSD are shown in Figure 7. AKD is a particle-based state entropy estimator used to evaluate skill activity, which we define in Section 4.4.
CIC is able to produce continuous movements of high activity, but it can’t generate behaviors at other activity levels (AKD range of CIC’s skills is 7-10). In addition, CIC suffers from insufficient exploitation, i.e., the generated skills are indistinguishable and homogeneous. CIC’s skills all tend to achieve dynamic flipping, which is consistent with the good initial score on walker flip in skill combination (see Section 4.2). However, the behavioral indistinguishability makes it difficult for meta-controllers to learn ideal combinations for a competitive final score. APS can generate diverse behaviors, but it suffers from lazy exploration and also can’t generate behaviors at different activity levels (AKD range of APS’s skills is 0-2). In general, poor exploration causes poor performance in skill finetuning evaluation. In skill combination, the diversity allows the meta-controller to complete downstream tasks through the combination of unqualified learned skills, which coincides with the upward trend of APS training curves (see Section 4.2). In summary, previous advanced methods can’t provide a good balance between behavioral quality and diversity, thus failing to exhibit competitive results across different downstream adaptation evaluations.
By contrast, our ComSD can produce diverse behaviors at different activity levels (AKD range of ComSD’s skills is 0-10), including flipping, lying down, struggling at different speeds, various postures, and so on. This explains why our ComSD can achieve state-of-the-art adaptation performance across both kinds of downstream tasks while other methods cannot.
D.2 Behavioral Diversity Analysis
In this section, we try to analyze the behavioral diversity of each method. For a skill, we use it to sample 1k states and calculate the Mean State (MS) over all sampled states. MS can partially represent the skill to some extent. For each method, we uniformly sample 41 different walker skills. Over 41 skills, we compute the K-Th-nearest-neighbor MS Distance (KTMSD) and the All-K-nearset-neighbor MS Distance (AKMSD) for skill entropy estimation. KTMSD computes the k-th-nearest-neighbor MS Euclidean distance of each skill and takes the mean, while AKMSD considers all the k-nearest neighbors of each skill and takes the mean distance of MS. (Note that in Section 4.4&Appendix D.1, KTD and AKD are employed for state entropy estimation of one skill, while KTMSD and AKMSD are used for skill entropy estimation of one method in this section.) KTMSD and AKMSD can partially evaluate the skill coverage of one method. In addition, we also calculate the MS Range (MSR) for each method. These metrics are all related to behavioral diversity.
@@ D.2 Behavioral Diversity Analysis
The comparison between all 6 methods on behavioral diversity is shown in Figure 8, demonstrating that ComSD has huge advantages on skill entropy (KTMSD and AKMSD). ComSD is also the most competitive method on MSR. In fact, it’s hard to represent an exploratory skill by only MS (exploratory behaviors have much more visited states than static postures), which puts highly exploratory methods (ComSD and CIC) at a disadvantage on these 3 metrics. In this case, ComSD still obtains state-of-the-art evaluation results, which demonstrates that the behavior set discovered by ComSD is of much higher diversity and coverage than other baselines.
@@
Figure 8: The comparison between all 6 methods on behavioral diversity. ComSD discovers a much more diverse skill set than other baselines.

### v2

@@ 5 Conclusion
In this paper, we propose ComSD to balance behavioral exploration and diversity through contrastive learning from a multi-objective perspective. We conduct extensive experiments and detailed analysis, verifying the superiority of ComSD on both skill quality and adaptation ability over all other baselines. We hope (i) our ComSD can inspire more attention to the exploration-diversity trade-off in complex behavior discovery, and (ii) our comprehensive experiments can inspire unsupervised skill evaluation for future works.
@@ A.3 Skill Finetuning Evaluation
Appendix B Extended Backgrounds
B.1 Related Works
Reinforcement Learning (RL) has proven its effectiveness in learning useful task-specific skills in the presence of extrinsic rewards  Mnih et al. [2015]; Levine et al. [2016]; Ding et al. [2021]; Narvekar et al. [2017]; Li et al. [2019]. The success of unsupervised learning in computer vision Chen et al. [2020]; Caron et al. [2020] and natural language processing Brown et al. [2020]; Devlin et al. [2018] further benefits task-specific RL with complex input Laskin et al. [2020]; Kostrikov et al. [2020]; Yarats et al. [2021a]; Stooke et al. [2021] by improving representation learning. However, the agents trained with lots of efforts are always hard to generalize their knowledge to novel tasks due to the task-specific supervision of extrinsic reward Stooke et al. [2021]. Unsupervised RL  Laskin et al. [2021] is proposed to improve this generalization, where the task-agnostic reward is designed for unsupervised pre-training. The pre-trained feature encoders  Liu and Abbeel [2021b]; Yarats et al. [2021b]; Liu et al. [2023] and exploration policies  Pathak et al. [2017, 2019]; Burda et al. [2018] then can be implemented for efficient RL on different downstream tasks. In addition to the generalization, intelligent agents should also be able to explore environments and learn different useful behaviors without any extrinsic supervision, like human beings. For the reasons above, unsupervised skill discovery  Gregor et al. [2016]; Eysenbach et al. [2018]; Park et al. [2021, 2023]; Strouse et al. [2021]; Shafiullah and Pinto [2022]; Kim et al. [2023]; Li et al. [2023]; Zhao et al. [2022]; Jiang et al. [2022]; Liu and Abbeel [2021a]; Lee et al. [2019]; Laskin et al. [2022b] is proposed and becomes a novel research hotspot. As a branch of unsupervised RL, unsupervised skill discovery also designs task-agnostic rewards and achieves unsupervised pre-training with these rewards, which guarantees task-agnostic downstream generalization. The main difference is that skill discovery requires extra input as conditions, named skill vectors. They aim to discover useful task-agnostic policies that are distinguishable by skill vectors. These approaches exhibit promising results across different fields  Brockman et al. [2016]; Bellemare et al. [2013]; Plappert et al. [2018]; Campos et al. [2020]. However, it’s hard for recent advanced methods to take both behavioral exploration and diversity into account, especially when faced with complex domains Laskin et al. [2021]; Tassa et al. [2018]. This is because their intrinsic rewards don’t provide effective incentives to seek and discern far-reaching skills for complex agents, and they also ignore the exploration-diversity conflict (see Section 3.3) when potential behaviors are hard to discern. By contrast, our ComSD designs strong incentives to respectively encourage diversity and exploration through contrastive learning, with a novel dynamic weight proposed for their balance.
B.2 Baseline Details
@@
Eysenbach et al. [2018] is one of the most classical and original unsupervised skill discovery algorithms, trying to maximize the MI between skills and states. It employs the first MI decomposition, Eq. (1). It uses a discrete uniform prior distribution to guarantee the maximization of skill entropy $H(z)$. The negative state-conditioned entropy $-H(z|s)$ is estimated by a trainable discriminator ${\rm log}p(z|s)$ which computes the intrinsic reward. As a foundational work, it provides several reasonable evaluations of skill adaptation, of which skill finetuning and skill combination are employed in our experiments.
@@
Lee et al. [2019] aims to learn a policy for which the state marginal distribution matches a given target state distribution. It optimizes the objective by reducing it to a two-player, zero-sum game between a state density model and a parametric policy. Like DIAYN, it is also based on the first decomposition (Eq. (1)) of MI and employs discriminator. The difference is that SMM explicitly maximizes state entropy with intrinsic reward, which inspires lots of recent works including ours.
@@
Liu and Abbeel [2021a] employs the second MI decomposition Eq. (2) for a better MI estimation. For state entropy estimation, it employs a popular particle-based entropy estimation proposed by APT Liu and Abbeel [2021b] for exploration, which is proven effective and supports many advanced works Laskin et al. [2022b]; Yarats et al. [2021b] and our ComSD. For skill-conditioned entropy, it chooses the successor feature Hansen et al. [2019] as a diversity reward with a fixed weight. However, we find that APS’s reward can’t provide enough incentives for complex agents, and they are also faced with the exploration-diversity conflict. To this end, our ComSD proposes to design strong incentives through contrastive learning and presents a novel dynamic weighting algorithm (SMW) to overcome the exploration drop brought by explicit conditioned entropy maximization (i.e., diversity reward).
@@
Laskin et al. [2022b] is a state-of-the-art robot behavior discovery method. It first introduces contrastive learning Chen et al. [2020] into unsupervised skill discovery. It chooses the second MI decomposition, Eq. 2 with APT particle-based estimation for state entropy, like APS. The contrastive learning between state transitions and skill vectors is conducted for implicit skill-conditioned entropy maximization. The encoder learned by contrastive learning is further used for APT reward improvement. The behaviors produced by CIC are of high activity but low diversity. Different from CIC, we employ the contrastive results as a novel diversity reward for explicit conditioned entropy maximization to improve behavioral diversity, with SMW for exploration maintenance.
@@
Yang et al. [2023] is another state-of-the-art method on URLB Laskin et al. [2021] and maze exploration Campos et al. [2020]. It tries to mitigate the exploitation problem in CIC by a novel MI objective, $I(s^{1},s^{2})$, where $s^{1}$ and $s^{2}$ denote different states generated by the same skill. It provides theoretical proof to show that their novel MI objective serves as the upper bound of the previous MI objective. However, BeCL can’t generate enough dynamic robot behaviors, and their intrinsic reward computational consumption is also much larger than other approaches.
@@ Appendix D Additional Analysis
D.1 Additional Diversity Analysis
In this section, we further analyze the behavioral diversity of each method. In Section 4.3, we use Mean State (MS) to represent one skill, and calculate the MS-coverage over several skills (selected by grid sweep Laskin et al. [2022b]) to represent the behavioral diversity of one method. In addition to MS-coverage, we further define another two metrics for behavioral diversity evaluation: MS-coverage-aknn and MS-range. Here we provide the concrete computation of these three metrics. MS-coverage: For one method, we calculate the MS Euclidean distance between each skill and its k-th-nearest neighbor skill, then take the mean. MS-coverage-aknn: For one method, we calculate the mean MS Euclidean distances between each skill and its all k-nearest neighbor skills, then further take the mean. MS-range: For one method, we report the maximum MS range of selected skills.
@@ D.1 Additional Diversity Analysis
The comparison between all 6 methods on behavioral diversity is shown in Figure 7. ComSD has huge advantages on two coverage metrics, demonstrating that its behaviors have the largest coverage in the potential skill space. Perhaps the most interesting phenomenon is that ComSD is also the most competitive method on MS-range. In fact, it’s hard to represent a dynamic behavior by MS (two different dynamic behaviors may have the same mean state), which puts highly exploratory methods (ComSD and CIC) at a disadvantage on these three metrics. In this case, CIC is at the bottom of all three metrics, while ComSD still obtains state-of-the-art evaluation results. This effectively demonstrates that ComSD produces behaviors of much higher diversity than all other baselines.
@@
Figure 7: The comparison between all 6 methods on 3 diversity metrics. ComSD discovers behaviors of much more diversity than other baselines.
D.2 Skill Visualization: What Skills Do ComSD and Competitive Baselines Discover?
Figure 8: Visualization for representative behaviors discovered by CIC and APS. AKD is a state entropy estimator used to evaluate skill exploration (activity), which we define in Section 4.3. Skill AKD ranges from 7 to 10 in CIC and 0 to 2 in APS, which means they can’t generate behaviors at different exploratory levels. In addition, CIC’s behaviors are also of low diversity.
Figure 9: Visualization for representative behaviors discovered by ComSD. AKD is a state entropy estimator used to evaluate skill exploration (activity), which we define in Section 4.3. Skill AKD ranges from 0 to 10 in ComSD. The results verify that our ComSD can produce diverse behaviors at different exploratory levels, which recent advanced methods (CIC and APS) cannot.
We provide the skill visualization and corresponding skill exploration (AKD) of the two most competitive baselines, APS Liu and Abbeel [2021a] and CIC Laskin et al. [2022b], in Figure 8. The visualization and skill AKD of our ComSD are shown in Figure 9. AKD is used to represent skill exploration (activity), which we define in Section 4.3.
CIC is able to produce continuous movements of high exploration, but it can’t generate behaviors at other activity levels (AKD range of CIC’s skills is 7-10). In addition, CIC suffers from insufficient diversity, i.e., the generated skills are indistinguishable and homogeneous. CIC’s skills all tend to achieve dynamic flipping, which is consistent with the good initial score on walker flip in skill combination (see Section 4.2). However, the low behavioral diversity makes it difficult for meta-controllers to learn ideal combinations for a competitive final score.
APS can generate diverse robot skills, but it suffers from lazy exploration and also can’t generate behaviors at different activity levels (AKD range of APS’s skills is 0-2). Compared with CIC, the higher diversity allows the meta-controller to partially complete downstream tasks, which coincides with the upward trend of APS training curves (see Section 4.2). However, the poor exploration causes poor performance of APS in skill finetuning, simultaneously.
In summary, previous advanced methods can’t provide a good balance between behavioral exploration and diversity, thus failing to exhibit competitive results across different downstream adaptation evaluations. By contrast, our ComSD can produce diverse behaviors at different exploratory levels (AKD range of ComSD’s skills is 0-10), including flipping, lying down, struggling at different speeds, various postures, and so on. ComSD’s huge superiority in behavioral quality explains why it achieves state-of-the-art adaptation performance across both kinds of downstream tasks, while other methods cannot.